OpenStudy (anonymous):

how do I figure out density with a volume of 10 cubic centimeters and a weight of 10 grams?

OpenStudy (anonymous):

Density is just the mass of something divided by its volume. It helps us to understand how "heavy" things are relative to one another. When I was at school, sugar cubes were exactly one centimter on each side and I used to imagine little cubes of everything. The metal cubes would be heaveir than the wooden ones and they would be heavier than the gaseous ones.

OpenStudy (anonymous):

Hi there, First of all think about the density. ρ = \[kg \div m^3\] this means that you need to divide the grams with the cubic centimeter. However, inorder to make the units consistent I would suggest to convert grams to Kg and \[cm^3 \to m^3\] To do this you need to think that one gram needs to be divided by a 1000 or 10^-3 to convert to Kg. Thus 10 grams = 0.01Kg Now convert \[cm^3 \to m^3\] Remeber to convert cm to meters you need to divide by 100. Since you have \[cm^3\] you need to divide by 3 times 100 or 10^-6. Finally, 10 cubic centimeters are 10 * 10^-6 cubic meters. Thus your answer is \[(10gr *10^-3)\div(10cm^3 * 10^-6) = 10^3 = 1000 kg/m^3\] and that is your density my friend wich is the density of water.